Operations Report – November 19th

Crew 215 Operations Report 19-Nov-2019
SOL: 9
Name of person filing report: Shane Usher
Non-nominal systems: N/A
Notes on non-nominal systems: N/A
Generator: run
Hours run: 12 h
From what time last night: 1910
To what time this morning: 0710
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 78%
Diesel Reading – 46%
Station Propane Reading – 67%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 42
Water Meter (units): 1459254
Water (static tank) (gal): 327
Static to Loft Pump used – yes
Water in Green Hab (gal): 243.7
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: Assigned to Director
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: yes
Hours: 111.7
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 41%
Currently charging: yes
Opportunity rover used: no
Hours: 68
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: yes
Hours: 118.1
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 63%
Currently charging: yes
Notes on rovers: Spirit had high drain rate during use with 2 passengers.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used (gal): 0
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With Director
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: Voice Activation Transmission had been activated on some radios and caused general confusion. VOX now deactivated on all radios.
Summary of Hab operations: N/A
Summary of GreenHab operations: Atmospheric equalization to the green hab carried out during the afternoon, plants were watered (20L,5.7 gal). Door closed at 5pm, heater on for a short time.
Summary of Science Dome operations: Gypsum cement experiments conducted.
Summary of RAM operations: N/A
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: Requested supplies arrived through RAM

Copyright © The Mars Society. All rights reserved. | Main Site